Circular No. 8891

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 17P/HOLMES M. Drahus and L. Paganini, Max-Planck-Institut für Sonnensystem-

forschung; and L. Ziurys, W. Peters, M. Soukup, and M. Begam, Arizona Radio Observatory (ARO), report their monitoring of comet 17P at millimeter wavelengths between Oct. 25.5 and 31.5 UT with the ARO 12-m telescope on Kitt Peak. Evolution of the outburst was traced through the HCN(1-0) transition. The following integrals of the antenna temperature were obtained, with the three hyperfine components summed, given in units of K km/s: Oct. 25.54, 5.0 ± 0.1; Oct. 27.58, 1.3 ± 0.1; Oct. 28.38, 0.58 ± 0.05; Oct. 30.40, 0.62 ± 0.05; Oct. 31.50, 0.39 ± 0.06. The gas-expansion velocity, measured on the blue wing of the strongest HCN line component, was fairly constant during the observations and equal to 0.50–0.55 km/s. The shapes and intensities of the HCN(1-0) triplet on Oct. 25.54 and 27.58 confirm the presence of the “chunk”, and the isotropic “bubble” (cf. CBETs 1111, 1118), the “chunk” being redshifted with respect to the nucleus. Both features were continually fading, likely due to photodissociation and because of moving away from the beam center. Since Oct. 28.38, the “chunk” has not been visible in the line shapes, and the intensity of HCN(1-0) has simultaneously stabilized; therefore, the measurements obtained since then likely result from a regular (non-explosive) post-outburst activity of the nucleus. On Oct. 26.51, five lines of CH_{3}OH were detected around 157 GHz, whose relative intensities give a rotational temperature of 50 ± 5 K. The area of the strongest line, 4(0,3)-4(1,4), was 0.78 ± 0.04 K km/s. Also detected were CS(3-2) on Oct. 26.58 (1.48 ± 0.04) and 29.38 (0.25 ± 0.02); H_{2}CO(2(0,2)-1(0,1)) on Oct. 27.53 (0.08 ± 0.04); and H_{2}S(1(1,0)-1(0,1)) on Oct. 31.41 (0.13 ± 0.02). Sensitive 3-sigma upper limits were obtained on CO(1-0) on Oct. 25.64 (< 0.2; however, a positive residuum of 0.05 is visible) and 27.46, (< 0.08). The half-power radius of the beam ranged from 19′′ to 36′′, depending on the observed frequency. The absolute calibration and stability of the instrument is estimated to be better than 20 percent.

COMETS C/2007 K21 AND C/2007 M10 (SOHO) Additional Kreutz-sungrazing comets (cf. IAUC 8890) — both of mag

∼ 7 and stellar, though C/2007 M10 appeared diffuse in C2 images:

Comet 2007 UT α_{2000} δ_{2000} Inst. F MPEC
C/2007 K21 May 23.358 3^{h}59.^{m}9 +18^{o}50′ C2  KB 2007-U23
C/2007 M10 June 29.846 6 23.8 +20 53 C3/2 BZ 2007-U16
2007 November 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green